# Instructions for ACL 2023 Proceedings

### v1

@@ A.7 Model Checkpoints Used
Appendix B Additional Experiments and Ablations
B.1 Captioning Qualitative Examples
We show in Figure 5 more examples of captions generated by captioning models trained with CA, T-MARS and ICC filtering. As can be seen, the model trained on the subset filtered with ICC generates more concrete and accurate captions, compared to CA and T-MARS filtering based model which, while generating captions which are somewhat related to the image, are highly abstract and less accurate. We use beam search of 5 for decoding in all models, including the examples shown in the main paper.
B.2 Qualitative examples Showcasing the Importance of Both Pipelines
@@
We visually show examples of each of the scores’ weaknesses and the way they compliment each other. In Figure 9, we show examples of concrete captions, the reconstructed captions by VBA and SBA, and the different scores of each of them. The first four rows exemplify why VBA may fail to reconstruct some concrete captions. For instance, the caption “a nurse mopping a surgeon’s brow during an operation in an operation pub” was reconstructed to “two people in protective gear” which bears relatively low semantic similarity to the original caption. The main reason these cases happen is due to the inherent difficulty of reconstructing (through a captioning model) from an image the exact caption from which the image was generated, as there may be many possible such captions. In this case, the use of SBA helps determining that the caption is concrete.
@@
In a complementary way, we show in Figure 10 examples of abstract captions. In this figure, the first four rows demonstrate that using SBA alone is also not enough, as it is sometimes able to reconstruct abstract captions due to the higher semantic information that is contained in the CLIP embeddings. In this scenario, VBA covers up for these failures, as it is very unlikely to reconstruct abstract text.
@@
These qualitative examples further illustrate the benefit of using both VBA and SBA. Indeed, in both Figure 9 and 10, it can be observed that ICC learns to take the best of both worlds, generating low scores for abstract captions, and high scores to concrete ones in a consistent manner.
Input caption SBA reconstructed caption VBA reconstructed caption VBA bottleneck image SBA VBA
ICC
a nurse mopping a surgeon’s brow during an operation in an operation pub
a nurse wiping the brow of a surgeon during an operation in an operating room
two people in protective gear
0.77
0.25
0.72
bougainvillea climbing up the wall of a villa
bougainvillea climbing on a wall of a villa
a house covered in pink flowers
0.72
0.26
0.81
table top shot of many vegetables and mexican bugs on a table
close up shot of vegetables and bugs on a table
vegetables arranged in the shape of a human head
0.70
0.25
0.76
silhouette of a man with a gun in poses royalty
silhouette of a man holding a gun in poses royalty
a group of people silhouettes on a white background
0.82
0.26
0.93
small flock of sheep in winter snow on a hilltop
small flock of sheep in snow on a hill
a herd of sheep in the snow
0.72
0.95
1.0
small blue and white airplane parked on the ramp with a control tower in the distance
small blue and white airplane parked on the tarmac next to a control tower
a blue and white airplane parked on the tarmac
0.96
0.95
1.0
a young girl runs through a field of cabbages
a young girl runs through a field of cabbages
a girl walking through a field of cabbage
0.96
0.95
1.0
a red post box and a telephone box stand together in a village
a red telephone box and a post box stand together in a village
a red post box next to a stone wall
0.84
0.89
0.92
@@
Input caption SBA reconstructed caption VBA reconstructed caption VBA bottleneck image SBA VBA
ICC
keep an eye on the ball when it comes to investments
keep an eye on the ball when it comes to investments
a soccer ball on a green field
0.91
0.19
0.1
what ’s the best thing about having a best friend of the opposite gender ?
the best thing about having a friend of the opposite gender
two young women sitting on a bench
0.89
0.16
0.1
film character : would you like to bet on these shares this christmas ?
which film character would you like to see in your shares this christmas?
santa claus, santa claus and sant
0.79
0.1
0
this is located in my home town !
this is located in my hometown!
a sign in front of a statue
0.75
0.28
0
chaotic systems are sometimes described using fractal patterns
fractals are patterns that can be found in many forms, such as chaotic systems and natural structures.
a black and white tunnel
0.22
0.19
0
on an average , the sloth travels feet a day
a sloth spends most of the day on its feet
a sloth hanging from a branch
0.17
0.27
0
get tips for biological genus , more commonly known as air plants , in your home
learn how to care for air plants, one of
a bunch of air plants on a brown surface
0.32
0.25
0
versatile and highly capable , there ’s more to this tiny camera than its giant zoom
this little camera packs a big punch with its zoom lens and 2
a camera on a wooden table
0.25
0.24
0
@@
a bundt wedding cake with white chocolate dripping, evergreens, pinecones and sugar powder to imitate snow | A young boy stands in front of a wall with height measurement marks and has his hand up to show | Elderly man with a cocktail during holidays | silver soda can and glass with white background - Stock Photo | foto of wallabies - Portrait of a wallaby in the nature - JPG | Brindle pendant in light grey with copper interior | Stock photo of homemade cookies and a cup of coffe | Girl in the gym lifting up the barbell - silhouette | Young blond woman wearing a dress in the forest | A fish eye photo of people climbing high ropes | Young alligators basking in the sunlight | Couple sat by basket full of grapes | vintage book and light bulb on wood table | tree with birds and birdcages vector image | A boy volunteer with birds on his shoulders | Colorful streamers hanging from the ceiling. | A glass bowl full of yellow cream and red and orange coloured fruit on pink & white background | A Chinese man walks past a billboard for a new commercial development which reads ’Shangrila is in your mind but | Shopping bags isolated on the white background | child tying his own shoes | A spotted harrier cruising to look for food | Young girl in a party dress looking bored and unhappy | Avatars of a male and in business suits. | a stack of miso chocolate chip cookies on a white plate | Bobcat in a hollow log
foto of florida-orange - Jacksonville skyline in orange background in editable vector file - JPG | Old man cleans tables at KFC to take home leftover food for family | c1913 to 1955 tall stack of music, opera and ballet books sg | Tropical Leaf Necklace 16 | welcome to india card with famous landmarks vector image | Set of empty picture frames for your own vector | castle hotel and spa wedding photos, ceremoy and reception | Radar monitor - Aircraft radar for airport with world map… | Grey and yellow consulting or planning concept infographics set | A darkened hall filled with server racks on either side and a silhouette of a Facebook worker at the end, | 100Ducati Desmoquattro at 2009 Seattle International Motorcycle Show 2.jpg | A film still shows two panels, with green ink. On one of them, the letters RELIC can just be made | spence cabin weddings | businessman in modern office writing ghostwriter in the air | large patio roof with adjustable louvres for outdoor seating weather protection and shading | pic of gesture - vector illustration of collection of hand gestures - JPG | chubby woman eating on scale stock photo, chubby woman on the scale eating a yogurt, isolate on white by iMarin | Fisherman on a small blue and white boat with fishing net between the waves of the sea. Liguria Italy | pictures of bedroom architectural details from hgtv | EHM water ionizer and alkaline water machine factory on sale | Christopher Boffoli’s photograph of a toy motorcycle rider, jumping over three toy cars and a slice of cheesecake | Pirate kids and their treasure | Incotex Benson Straight Leg Wool Trousers | QR Code for Florida Virtual School at local Shell Gas Station | Peace Love Colorectal Surgery Oval Sticker (50 pk) | moonstruck chocolates | creative eyelashes - closeup of the eye of woman with… | Simply Perfect Braided Wedges | New Years Fireworks in Seattle, 2011->2012 | 2008 Volkswagen GTI Photo | A sport fishing boat heading out of Wanchese harbor of the Outer Banks at dawn for a day of off-shore | A young beautiful girl holding a wild fox animal that was traumatized by a man and rescued by her and
Zero Zebra Safari Party Dairy-Free Chocolate Animals | The rescued soccer team members pose with a sketch of the Thai Navy SEAL diver who died while trying to | How to get a bobcat out of your window blinds | Pocket watch: technically interesting pocket watch with rare crown winding in manner of O. | I am an Aspie Girl A Book for Young Girls with Autism Spectrum Conditions by Danuta Bulhak-Paterson, Tony Attwood | Trump on etch a sketch | floor plan drawing software create your own home design easily | 1000 ideas about lake house plans on pinterest house for Basement planner online | Chanel 5, the first perfume i received as a gift. Love it! | lush greenery at National gallery of modern arts - Bangalore | A lounge room of greys and creams, black and white prints all come together to make this a relaxing and | christmas bible verses for preschoolers five scriptures about children should 664 | MAJESTIC PET PRODUCTS - Santorini Chevron Round Pet Bed - pet bed looks great in any room of your house | Sanusi under house arrest, moved to a 2-bedroom apartment without electricity & access roads (Photos | Novak Djokovic (Ser) defeated Juan Martin Del Potro (Arg) in US Open final<br /> Flushing Meadows 09-09-2018 US Open<br /> | 2013 men’s the novelty original t-shirt with patterns Double-headed eagle and RUSSIA sizel xl xxl xxxl 4xl shirts free shipping | Poster of Seven Below | EFCC operatives evacuating the safe, the house where the cash was hidden and the money | closet converted into mudroom | make a closet more functional by removing doors, adding a bench at kid height, hooks | How to install an SSD in a laptop | computer tutorial | The logo of German carmaker BMW is seen on a car displayed during the annual results press conference in Munich, | Well Established and Well Equipped Butchers, Fruit and Vegetables, Frozen Seafood Plus Convenience Groceries, South London for sale | BMW Is Looking Into Gas-Powered Vehicles | summer fashion scarvesnew scarf trendswhite by scarvesCHIC on Etsy, \$15.90 | Celebrating 20 years, EGEC shares declaration on the great role of geothermal energy | capricorn tattoos designs ideas and meaning tattoos for you | 25 best ideas about spice storage on spice | Julbo Eyewear - Atmo Goggle (4-8 Years Old) (Red Trans Orange Lens) Snow Goggles | The Roman temple of Jupiter is seen in the background as Lebanese youths play in the snow on January 9, | The DJI Osmo kit includes the grip, gimbal, camera and device holder for your smartphone (there is a companion app). | Miyake celebrates with his team after winning a silver medal at London 2012. | National MS Society’s Katie Boothroyd, Board Member Joan Ohayon. Photo by Tony Powell. Tea Honoring Women of the Diplomatic Corps. | Top sweet and fortified wines of 2012 | there is also a small laundry with all-in-one washer and dryer | change my background how to change desktop background in windows 10 | Hand wrapping Basics - How to wrap your hands for boxing, kickboxing, and Muay Thai with long wraps | favorite colors on taupe benjamin and paint colors | Black & White Houndstooth Infinity Pocket Scarf - Travel Scarf - The Poppy Stock | looking out to the yew garden | An entrance door to transform your home for Home front entry doors | Image result for paytm with modi advertisement | Drawing notebook. never thought of including this inside a felt book, always had a separate art bag.. | how to remove a kitchen tile backsplash | ONLY - Cara Long Sleeve Shirt (Navy Blazer) Women | Tying an olive dun with mallard wings | All cold and hot rolling seamless steel pipe diameter | and when the clock strikes midnight halloween will be over | Another member said she’d tried for years to get her floor to look this clean, declaring she ’loved’ the bargain | 2PCS/Set Hematite Natural Stone Bracelet1 | architectural drawings for sale unknown vintage architectural drawing for sale at 1stdibs | Prince Harry to lose honorary military titles as palace confirms exit date from royal family | Explore One 70mm Telescope | The mess I left at the end of the school year in 2015 when I couldn’t unpack or move into | A pretty paper garland adds a festive touch. | Hydroponic Fodder ProFeed Growing System | Antioxidant rich RED salad with lemony dressing is delicious lunch! | Sunflowers on Saturday, when I felt called to ask David to photograph me with them since they played such a | pencil crafts for back to school and beyond | COLROVIE Culotte Leg Elegant Cami Jumpsuit Women Box Pleated Sexy V Neck Jumpsuits 2017 Fall Surplice | how to make fabric flower rosettes, tutorial | new cars with best warranty all about extended auto warranty contracts leadhub | 30 sliding barn door designs and ideas for the home With barn door wide opening | Foreign stocks for students and grannies | how to wear flat shoes to a wedding | A cartoon on the situation with languages in Ukraine cartoon. | Modest partisan differences in views of elected officials | interior design for my home minimalist interior design is maximum on style | Ebook download and read online electronic book button or icon | what do you want to know about alta motors electric street tracker | Raw Oysters are the perfect food to increase your testosterone! | Minecraft cake - Both tiers are vanilla cake with vanilla buttercream covered in fondant. The tiles are all modeling chocolate | cbd infographic why patients are leaving big pharma | the majestic elephant - one of the big 5 |
Car insurance advice: How to keep your car safe in winter weather | We only supply the tire. If there is a rim shown in the picture, it is for display purposes only. | "Steven Wright Quote: ""It’s like the Wild West, the Internet. There are no rules.""" | pathandpuddle: How long do animals live? | A lack of sleep could be caused the nutrients in your diet | No time to explain. Just put on the hats and act casual. | His last at-bat, a pop fly to center field. #garrettreade #littleleague #thatsmyboy | Higher consumption of sugary beverages linked with increased risk of mortality | Words of Gymnastics Terminology w/ Monogram Drawstring Bag | Zaanse Schans, Netherlands - May 5, 2015: Tourist Visit Windmills And Rural Houses In Zaanse Schans | Losing out: BP will temporarily be locked out of lucrative deals, including contracts to supply the US military with fuel. | QuickBooks - Access | "Cranberry Chevron Rug - Deep red hues cut a rug here. The chevron is a ""go with anything"" pattern and | Bomag reports that their single direction vibratory plate compactors are great tools for contractors’ day-to-day use in soil and asphalt | "Augustabernard bias-cut satin evening gown, c.1930. Label: ""Augustabernard"" with a stamped couture number on the back." | How to graduate as a successful edupreneur | garland for stairs christmas house tour decorating ideas how decorate for | This Mexican Layer Dip is easy to make and full of flavor! With layers of spicy black bean dip, homemade | what is a research process paper The term research paper may also refer to a scholarly article that contains the | Ammonia is often used in cleaning products because it reacts with grease making this easier to remove | They do not take ownership of valuable deposited with them? | Can i push out my wall to get an 8x8 bathroom leave me for Small bathroom design 5 x 8 | Making one of these wall hangings is a great way to use up old yarn … | 6 Pack - 16 Ounce Grolsch Bottles with Easy Cap Flip Top Caps for Brewing Beer, Kombucha, Kefir, Water, Thick | Annual: The event, celebrated every year to herald good monsoon rains for increased rice harvest, prosperity and goodluck, is one | Graphic on Australia’s Tasmanian Devils, rare carnivorous marsupials in a battle for survival against a contagious facial cancer. It’s been | remodeling small bathroom ideas on a budget small bathroom remodel on a budget brown ceramic tile floor walk | When a 17-year-old Sharapova burst onto the scene in 2004 by winning Wimbledon, she was an immediate hit. Sure, her | The course has a convenient location in the community of Fernie. | Famous burj al arab hotel dubai 6 said to be the for Dubai world famous hotel | Among Us. (Innersloth) | Download car and vehicles decal graphics kit designs ready to print and cut for vinyl | GSM Functionality GSM Technology is a special design which can be used in conjunction with a variety of signallers depending | Kendal Town Council are calling for a bypass to solve traffic problems after Storm Desmond | Are You Am I - Faira Dress **white | interior home design also with a interior wall design also with a | One reason Millennials book cruises is the low cost - contrary to popular belief | Containers have grabbed a large share of the intermodal freight business, and here’s a miniature trainload of them at the | be a donor be a hero | "Operation Surf Santa Cruz is an annual event that honors active duty military soldiers through ""an epic life-changing surfing experience.""Many | hyster h40 h forklift will not go forward or backward stuck brakes | Nominated banks roles and responsibilities under a letter of credit transaction. | Kingsbrook animal hospitals blog preferred veterinary care in two beloved basenjis kylie and cricket in a house fire in april | Example of TRAM:Cross-Domain-Solution | Skip the recycling. Use your soda cans to make bracelet cuffs instead! | They rise in quite an interesting way as well. Keep an eye on them so they don’t burn, and make | "15 Thanksgiving Day Ideas for Couples. Holidays get so busy we sometimes forget our priorities. Keep your spouse at the | Tao Xiangli gets out of his homemade submarine after operating it in a lake on the outskirts of Beijing September | Professionally edit your voiceover, audiobook or podcast | "Eli struggling to find cell service ""under the biggest cell phone tower in Paris""" | Don’t have time to shampoo your hair, but still want to look glam? Me too! And that’s why I love | Vivid Vision for Success | Rich Karlgaard | The pistol 01 concept is a very precise and reliable sidearm! | buy a domain name | Tate Stevens - Winner of 2012 X Factor, Simon Cowell Stock Photo | Click this cover for a(n) eBook sample of Choke Point | FI Week #5 #6 #7 : Let the new digital influencers shine | Travertine is a perfect partner2 | STF has arrested the wanted accused who recruited fake teachers | SUMMER HITS 2016 Mixed by DJ Golan | Springet fra Big Business til Smaller Business (Starting Over) | Banque De Francs, French Banknote Assortment, ca. 1945-1961. | tarek christina tarek and christina el moussa s divorce affects their | Phoenix Suns are changing their perception to the basketball world | Tan Cartoon Doctor Man Carrying His - African American… | 2013 New arribal digital holly quran mp4 player wholesale price and 28 language translation(China (Mainland)) | Your spa decor should reflect your target market’s preferences. | hwaseong-fortress-suwon-part-2 | Salman and Sonakshi in Dabangg 2 |
Figure 11: Manually Annotated Captions. The captions are sorted according to concreteness, where captions with the highest score illustrated in the top cluster and lowest at the bottom cluster. We truncate captions that are longer than 20 words, and separate captions by ||||.

### v4

@@ A.8 Model Checkpoints Used
A.9 Finding the Score Combination Parameters
To compute the combination parameters of the SBA and VBA scores, we label 244 captions, sampled uniformly over VBA and SBA scores, with concreteness scores in the range 003333. We use logistic regression to find the parameters $a,b,c$ of $\sigma(a\cdot VBA{}+b\cdot SBA{}+c)$, where $\sigma(x)=\frac{1}{1+e^{-x}}$ is the sigmoid function, such that the output will approach 1 for concrete captions and 0 for abstract ones. We label concrete captions as captions with concreteness above the median score in the labeled dataset and abstract captions as captions with a score below this median. We visualize the annotated samples and the regression line $a\cdot VBA{}+b\cdot SBA{}+c=0$ in Figure 6. The parameters found and used in our ICC are $a=13.2$, $b=3.6$ and $c=-9.4$. As seen in the figure, both scores contribute to the optimal predicted concreteness score, validating the importance of using both SBA and VBA components together in our full pipeline.
Appendix B Additional Qualitative Examples
@@
We visually show examples of each of the scores’ weaknesses and the way they compliment each other. In Figure 8, we show examples of concrete captions, the reconstructed captions by VBA and SBA, and the different scores of each of them. The first four rows exemplify why VBA may fail to reconstruct some concrete captions. For instance, the caption “a nurse mopping a surgeon’s brow during an operation in an operation pub” was reconstructed to “two people in protective gear” which bears relatively low semantic similarity to the original caption. These cases mainly stem from the inherent difficulty of reconstructing (through a captioning model) from an image the exact caption from which the image was generated, as there may be many possible such captions. In this case, the use of SBA helps determining that the caption is concrete.
@@
In a complementary manner, we show in Figure 9 examples of abstract captions. In this figure, the first four rows demonstrate that using SBA alone is also not enough, as it is sometimes able to reconstruct abstract captions due to the higher semantic information that is contained in the CLIP embeddings. In this scenario, VBA compensates for these failures, as it is very unlikely to reconstruct abstract text.
@@
These qualitative examples further illustrate the benefit of using both VBA and SBA. Indeed, in Figures 89, it can be observed that ICC reflects the advantages of both pipelines by generating low scores for abstract captions, and high scores for concrete ones in a consistent manner.
Input caption SBA reconstructed caption VBA reconstructed caption VBA bottleneck image SBA VBA ICC a nurse mopping a surgeon’s brow during an operation in an operation pub a nurse wiping the brow of a surgeon during an operation in an operating room two people in protective gear Refer to caption 0.77 0.25 0.72 bougainvillea climbing up the wall of a villa bougainvillea climbing on a wall of a villa a house covered in pink flowers Refer to caption 0.72 0.26 0.81 table top shot of many vegetables and mexican bugs on a table close up shot of vegetables and bugs on a table vegetables arranged in the shape of a human head Refer to caption 0.70 0.25 0.76 silhouette of a man with a gun in poses royalty silhouette of a man holding a gun in poses royalty a group of people silhouettes on a white background Refer to caption 0.82 0.26 0.93 small flock of sheep in winter snow on a hilltop small flock of sheep in snow on a hill a herd of sheep in the snow Refer to caption 0.72 0.95 1.0 small blue and white airplane parked on the ramp with a control tower in the distance small blue and white airplane parked on the tarmac next to a control tower a blue and white airplane parked on the tarmac Refer to caption 0.96 0.95 1.0 a young girl runs through a field of cabbages a young girl runs through a field of cabbages a girl walking through a field of cabbage Refer to caption 0.96 0.95 1.0 a red post box and a telephone box stand together in a village a red telephone box and a post box stand together in a village a red post box next to a stone wall Refer to caption 0.84 0.89 0.92
@@
Input caption SBA reconstructed caption VBA reconstructed caption VBA bottleneck image SBA VBA ICC keep an eye on the ball when it comes to investments keep an eye on the ball when it comes to investments a soccer ball on a green field Refer to caption 0.91 0.19 0.1 what ’s the best thing about having a best friend of the opposite gender ? the best thing about having a friend of the opposite gender two young women sitting on a bench Refer to caption 0.89 0.16 0.1 film character : would you like to bet on these shares this christmas ? which film character would you like to see in your shares this christmas? santa claus, santa claus and sant Refer to caption 0.79 0.1 0 this is located in my home town ! this is located in my hometown! a sign in front of a statue Refer to caption 0.75 0.28 0 chaotic systems are sometimes described using fractal patterns fractals are patterns that can be found in many forms, such as chaotic systems and natural structures. a black and white tunnel Refer to caption 0.22 0.19 0 on an average , the sloth travels feet a day a sloth spends most of the day on its feet a sloth hanging from a branch Refer to caption 0.17 0.27 0 get tips for biological genus , more commonly known as air plants , in your home learn how to care for air plants, one of a bunch of air plants on a brown surface Refer to caption 0.32 0.25 0 versatile and highly capable , there ’s more to this tiny camera than its giant zoom this little camera packs a big punch with its zoom lens and 2 a camera on a wooden table Refer to caption 0.25 0.24 0